### v2

@@ 5.1 Random Evaluation
Since $I-X$ has ones along the diagonal and every other entry has constant term zero, we know that $\det(I-X)$ is a polynomial with constant term $1$, so by Proposition 3.1 the multiplicative inverse of $\det(I-X)$ is a well-defined formal power series. Thus, Equation 12 can be viewed as an equality between two matrices of formal power series.
@@
of Equation 11 over $\mathbb{F}$ is nonzero with probability at least $1-2nm/(2^{q})$.
Since we picked $q$ such that $2^{q}\geq 2n^{6}\geq n^{3}\cdot(2nm)$, the desired result follows.
@@ 6 Bounded Connectivity
    computes ranks of submatrices, whose dimensions are based off degrees of nodes in $G$.
@@ 6.1 Degree Reduction
Fix $s,t\in V$. Given an $st$-path $P^{\prime}$ in $G_{\text{new}}$, we recover a unique $st$-path $P$ in $G$ by looking at the sequence of original vertices $P^{\prime}$ passes through. Using this construction, any collection of $r$ edge-disjoint paths from $s$ to $t$ in $G_{\text{new}}$ recovers a collection of $r$ edge-disjoint paths from $s$ to $t$ in $G$. So the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\lambda(s,t)$.
@@
Thus the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\min(k,\lambda(s,t))$.
Set $\lambda=\min(k,\lambda(s,t))$.
@@ 6.2 Low-Rank Edge-Adjacency
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\omega(\mathcal{C})\right).$ (17)
For each collection of walks $\mathcal{C}$, let $\tilde{\omega}(\mathcal{C})$ be the monomial resulting from substituting Equation 16 into the weight $\omega(\mathcal{C})$. Then we have
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\tilde{\omega}(\mathcal{C})\right).$ (18)
@@
$\tilde{\omega}(\mathcal{P})=\prod_{i=1}^{r}\prod_{(e,f)\in\mathcal{E}_{i}}\left(\sum_{j=1}^{k}y_{ej}z_{jf}\right).$
@@
Figure 3: Given edge-disjoint paths $P_{1}=\langle e_{1},f_{1}\rangle$ and $P_{2}=\langle e_{2},f_{2}\rangle$ in $G$, the determinant of the matrix $(I-X)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ enumerates this pair via the monomial $\omega(P_{1},P_{2})=x_{e_{1}f_{1}}\cdot x_{e_{2}f_{2}}$. The variables in this monomial provide enough information to uniquely recover $P_{1}$ and $P_{2}$. In contrast, for $k=2$, the determinant of $(I-YZ)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ assigns this pair weight $\tilde{\omega}(P_{1},P_{2})=(y_{e_{1}1}z_{1f_{1}}+y_{e_{1}2}z_{2f_{1}})(y_{e_{2}1}z_{1f_{2}}+y_{e_{2}2}z_{2f_{2}}).$ One of the terms in the expansion of the above product is $y_{e_{1}1}z_{1f_{1}}\cdot y_{e_{2}2}z_{2f_{2}}$. We can read this term as saying “the first path $P_{1}$ traverses $e_{1}$ and $f_{1}$, and the second path $P_{2}$ traverses $e_{2}$ and $f_{2}$.” So this monomial provides enough information to recover the pair of paths $\langle P_{1},P_{2}\rangle$ as well.
@@ 6.4 The Algorithm
Step 2 of Algorithm 2 involves computing ranks of $n^{2}$ separate $k\times k$ matrices, which by Proposition 2.2 takes $\tilde{O}(k^{2}n^{\omega})$ time.
@@ 7 Conclusion
The general approach of solving graph problems by testing whether certain enumerating polynomials are nonzero is now a common technique in graph algorithms: we refer the reader to [EKW23] and the citations therein for previous examples of this paradigm in the literature.
@@
In the 4-Clique problem, we are given an undirected graph $G$ on $n$ nodes, and are tasked with determining if $G$ contains four vertices which are mutually adjacent. The best conditional lower bounds for APC come from observing that 4-Clique reduces to APC [AGI+18, Section 4]. Using the current best algorithms for rectangular matrix multiplication [WXXZ23, Table 1], the fastest known algorithm for 4-Clique takes $O(n^{3.251})$ time [EG04, Theorem 1]. Even if we conjecture this runtime time is optimal, the resulting lower bound for APC is a far cry from the $n^{4+o(1)}$ runtime we have for APC in dense graphs. Moreover, if $\omega=2$ this lower bound for APC weakens to $n^{3-o(1)}$.
@@
In undirected graphs, although APC can be solved in $\tilde{O}(n^{2})$ time, APVC is not known to admit a near-quadratic time algorithm. In fact, the best known lower bounds for APC in directed graphs also hold for APVC in undirected graphs [HLSW22].

### v3

@@ 5.1 Random Evaluation
Since $(I-X)$ has ones along the diagonal and its other entries have constant term zero, we know that $\det(I-X)$ is a polynomial with constant term $1$, so by Proposition 3.1 the multiplicative inverse of $\det(I-X)$ is a well-defined formal power series. Thus, Equation 12 can be viewed as an equality between two matrices of formal power series.
@@
of Equation 11 over $\mathbb{F}$ is nonzero with probability at least $1-2nm/2^{q}$.
Since we picked $q$ such that $2^{q}\geq 12n^{6}>n^{3}\cdot(2nm)$, the desired result follows.
@@ 6 Bounded Connectivity
    computes ranks of submatrices, whose dimensions depend on degrees of nodes in $G$.
@@ 6.1 Degree Reduction
Fix $s,t\in V$. Given an $st$-path $P^{\prime}$ in $G_{\text{new}}$, we recover a unique $st$-path $P$ in $G$ by looking at the sequence of original vertices visited by $P^{\prime}$. Using this construction, any collection of $r$ edge-disjoint paths from $s$ to $t$ in $G_{\text{new}}$ recovers a collection of $r$ edge-disjoint paths from $s$ to $t$ in $G$. So the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\lambda(s,t)$.
@@
Thus the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\min(k,\lambda(s,t))$. Set $\lambda=\min(k,\lambda(s,t))$.
@@ 6.2 Low-Rank Edge-Adjacency
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$ (17)
For each collection of walks $\mathcal{C}$, let $\tilde{\xi}(\mathcal{C})$ be the monomial resulting from substituting Equation 16 into the weight $\xi(\mathcal{C})$. Then we have
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\tilde{\xi}(\mathcal{C})\right).$ (18)
@@
$\tilde{\xi}(\mathcal{P})=\prod_{i=1}^{r}\prod_{(e,f)\in\mathcal{E}_{i}}\left(\sum_{j=1}^{k}y_{ej}z_{jf}\right).$
@@
Figure 3: Given edge-disjoint paths $P_{1}=\langle e_{1},f_{1}\rangle$ and $P_{2}=\langle e_{2},f_{2}\rangle$ in $G$, the determinant of the matrix $(I-X)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ enumerates this pair via the monomial $\xi(P_{1},P_{2})=x_{e_{1}f_{1}}\cdot x_{e_{2}f_{2}}$. The variables in this monomial provide enough information to uniquely recover $P_{1}$ and $P_{2}$. In contrast, for $k=2$, the determinant of $(I-YZ)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ assigns this pair weight $\tilde{\xi}(P_{1},P_{2})=(y_{e_{1}1}z_{1f_{1}}+y_{e_{1}2}z_{2f_{1}})(y_{e_{2}1}z_{1f_{2}}+y_{e_{2}2}z_{2f_{2}}).$ One of the terms in the expansion of the above product is $y_{e_{1}1}z_{1f_{1}}\cdot y_{e_{2}2}z_{2f_{2}}$. We can read this term as saying “the first path $P_{1}$ traverses $e_{1}$ and $f_{1}$, and the second path $P_{2}$ traverses $e_{2}$ and $f_{2}$.” So this monomial provides enough information to recover the pair of paths $\langle P_{1},P_{2}\rangle$ as well.
@@ 6.4 The Algorithm
Step 2 of Algorithm 2 involves computing ranks of $n^{2}$ separate $k\times k$ matrices, which by Proposition 2.2 takes $\tilde{O}(k^{\omega}n^{2})$ time.
@@ 7 Conclusion
The general approach of solving graph problems by testing whether certain enumerating polynomials are nonzero is now a common technique in graph algorithms: we refer the reader to [EKW24] and the citations therein for previous examples of this paradigm in the literature.
@@
In the 4-Clique problem, we are given an undirected graph $G$ on $n$ nodes, and are tasked with determining if $G$ contains four vertices which are mutually adjacent. The best conditional lower bounds for APC come from observing that 4-Clique reduces to APC [AGI+19, Section 4]. Using the current best algorithms for rectangular matrix multiplication [WXXZ23, Table 1], the fastest known algorithm for 4-Clique takes $O(n^{3.251})$ time [EG04, Theorem 1]. Even if we conjecture this runtime time is optimal, the resulting lower bound for APC is a far cry from the $n^{4+o(1)}$ runtime we have for APC in dense graphs. Moreover, if $\omega=2$ this lower bound for APC weakens to $n^{3-o(1)}$.
@@
In undirected graphs, although APC can be solved in $\tilde{O}(n^{2})$ time, APVC is not known to admit a near-quadratic time algorithm. In fact, the best known lower bounds for APC in directed graphs also hold for APVC in undirected graphs [HLSW23].
@@
Derandomization
The algorithms we presented for APC and $k$-APC in this work are randomized. Can we solve these problems with respective $\tilde{O}(m^{\omega})$ and $\tilde{O}((kn)^{\omega})$ runtimes deterministically?